What must be the diameter of a hose if it is to deliver 8 liters of oil in 0.6 minutes with an exit velocity of 6 m/s?
Answer #1

here,

the volume of flow rate , V = 8/0.6 litre/min

V = 8/0.6 * (10^-3 /60) m^3 /s

V = 2.22 * 10^-4 m^3 /s

the exit velocity , v = 6 m/s

let the diameter of hose be d

V = Area * v

V = pi * (d/2)^2 * v

2.22 * 10^-4 = pi * ( d/2)^2 * 6

solving for d

d = 6.87 * 10^-3 m

the diameter of hose is 6.87 * 10^-3 m